OpenStudy (anonymous):

5x+3y=19 -5x+2y=-4 i will you addition and the solution is ( , ) HELP!

OpenStudy (anonymous):

5x +3y = 19 -5x+2y= -4 5y = 15 5y/5 =15/5 15/5 =3 y=3

OpenStudy (anonymous):

plug 3 back into 1st equation,5x+3(3)=19 5x+9 = 19 5x+9-9=19-9 5x=10 5x/5 =10/5 x=2

OpenStudy (anonymous):

thank you so much!